# Long-Term Fairness Inquiries and Pursuits in Machine Learning: A Survey of Notions, Methods, and Challenges

### v2

@@ 4.3 Discussion
Compared with the standard static fairness interventions, the summarized mitigation methods mainly contribute to formulating the dynamics of human-model interactions and establishing the corresponding solutions based on the group utility shifts. Long-term fairness under unknown dynamics and/or without leveraging demographic information remains largely unexplored but promising, especially considering privacy laws (Chen et al., 2019a) that limit access to such data. From the perspective of temporal learning, it is positive that long-term fairness is focused on the multi-step sequential learning setting. However, the current research typically only focuses on optimizing the model learner at each time step. Understanding the tensions between ensuring fairness at each time step and cumulatively at the final step needs more investigation. Finally, while RL methods are flexible and learn the dynamics of the environment well, they are data-intensive (Dulac-Arnold et al., 2021; Henderson et al., 2018) and often unsuitable for real-world scenarios. Hence, developing techniques that can efficiently estimate the dynamics is crucial. Notably, the recent advance in performative prediction Perdomo et al. (2020) provides a new perspective on off-policy learning for modeling the interplay between the model policy and agents.
@@ 5 Evaluation of Long-Term Fairness
Practical applications offer a unique context for long-term fairness considerations, depending on which one would expect context-dependent interpretations of technical findings. Therefore in this section, we present application-specific fairness inquiries motivated by empirical scenarios.
@@ 5.1 Application-Specific Studies
Credit Application. Liu et al. (2018) focus on lending practices to quantify delayed impacts of decisions (selections) made by institutes (e.g., the bank) on different groups. They represent each group through a score distribution, where higher scores indicate a greater likelihood of positive outcomes. Assuming access to a function that provides the expected score change for individuals, they introduce a one-step feedback model and construct outcome curves, which describe the relation between the expected difference in the mean score after the one-step feedback and the natural parameter regimes (e.g., selection rate of the policy). They found that both Demographic Parity and Equal Opportunity can lead to all possible outcomes, i.e., long-term improvement, stagnation, and decline, in the natural parameter regimes. This study emphasizes the impossibility of foreseeing a fairness criterion’s impact without carefully modeling the delayed outcome.
@@
Labor Market. The seminal work of (Hu & Chen, 2018) considers a two-stage dynamic reputational labor market model to capture the long-term and reinforcing nature of disparate outcomes. They propose imposing demographic parity within a temporary labor market to achieve long-term equilibrium in the permanent reputational labor market based on affirmative action. Separately, Celis et al. (2021) investigate the long-term impact of the Rooney Rule (Collins, 2007), a hiring policy aimed at mitigating implicit bias by mandating decision-makers to include at least one candidate from an underrepresented group. They employ a mathematical model of implicit bias, featuring two candidate groups, one of which is underrepresented. Each candidate possesses both a true, latent utility and an observed utility, subject to bias by a multiplicative factor. Candidates are then shortlisted by a selection panel based on observed utility. Their findings demonstrate that the rule progressively enhances minority hiring while maintaining overall utility.
@@
Predictive Policing Ensign et al. (2018) explore resource allocation in predictive policing (Mohler et al., 2015), which deploys patrol officers based on historical crime data. They identify a feedback phenomenon present in similar algorithms across contexts like recidivism prediction and lending, where prediction outcomes heavily influence algorithm feedback. Using the Pólya urn model, they explain this feedback occurrence and propose black-box solutions, such as reweighing input incidents. Elzayn et al. (2019), go one step further by proposing an algorithm which, under simplification assumptions, converges to an optimally fair resource allocation, even in the absence of known stationary ground-truth crime rates, avoiding the runaway feedback loop effect.
@@ 5.2 Long-Term Fairness in Recommender Systems
Long-term fairness considerations naturally align with Recommender Systems due to evolving user interactions and item popularity, impacting system performance Liu et al. (2021). Consequently, the RL framework is favored for its ability to model interactive dynamics. The works mostly focus on 1) user-side fairness, addressing equitable treatment for users, and 2) item-side fairness, ensuring unbiased item recommendation stemming from popularity bias due to disproportionate exposure Chen et al. (2023).
@@
Long-term fairness efforts in recommendation systems primarily target item-side fairness, addressing the challenge by modeling the sequential interactions between consumers and recommender systems using an MDP framework. A few works in this direction adopt an Actor-Critic RL algorithm leveraging the policy-based method (actor) with the value-based approach (critic) to learn the policy and value-function simultaneously Grondman et al. (2012). In this setup, the actor learns an optimal policy through environment exploration, while the critic evaluates actions to determine long-term rewards Liu et al. (2021). Utilizing this, Liu et al. (2021) propose FairRec that dynamically maintains accuracy and user fairness in the long term by jointly representing the user preferences (at the item and group level) as the states in an MDP model. They use the Weighted-Proportional Fairness concept, allocating resources (recommendations) proportionally to diverse entities’ demands using a weighting factor. This dual-reward system strategically optimizes both accuracy and fairness in the long term. Separately, Ge et al. (2021) study the long-term fairness of item exposure, focusing on group item labels. They address the challenge of items categorized into groups based on popularity, which dynamically change with user interaction over time, leading to shifting group labels. Using a Demographic Parity-based fairness-constrained policy optimization method, the authors propose an MDP recommendation model based on the actor-critic architecture. However, they introduce an additional critic designed for the demographic parity fairness constraint. On a related note, Morik et al. (2020) explored the rich-get-richer phenomenon in dynamic Learning-to-Rank approaches Yang & Ai (2021). Their solution, FairCo, ensures meritocratic fairness by allocating exposure to items proportional to their average relevance. FairCo integrates user feedback and dynamically adjusts its ranking based on the fairness constraint, employing a P-Controller approach for modeling, a popular feedback control mechanism (Bequette, 2003). In another work, Saito et al. (2020) tackles self-selection bias (Chen et al., 2023), where popular items tend to receive more clicks regardless of actual interest. They propose an ideal loss function optimized for relevance rather than the highest click probability. In this regard, uniform data has also been proven to be an effective approach Akpinar et al. (2022). This method involves randomly selecting items and ranking them using a uniform distribution. Chen et al. (2023) (Section 4.8) summarizes the various methods that use this approach. Lastly, Ferraro et al. (2021) investigate gender fairness in collaborative filtering, revealing disparities in exposure between female and male artists. They introduce an iterative re-ranking method that simulates feedback loops and penalizes male artists, resulting in adjusted rankings without significant performance impact.
@@
Relatively fewer works have focused on long-term user-side fairness. Akpinar et al. (2022) employ a simulation framework and Pólya urns to study the long-term impacts of fairness interventions in social network connection recommendations. They argue that static fairness notions fail to maintain equity in network sizes and utility across different protected groups in dynamic settings.
@@ 5.3 Discussion
The majority of the long-term fairness evaluations have either been context-specific (e.g., hiring) or in recommendation systems. In recommendation systems, RL methods are favored for modeling dynamics. However, a notable drawback of RL methods is their vulnerability to user tampering (Kasirzadeh & Evans, 2023), where recommender systems manipulate user opinions and preferences to boost long-term user engagement through its recommendations. Furthermore, assessing long-term fairness in a multi-stakeholder (Abdollahpouri & Burke, 2019) setting characterized by competing interests remains largely unexplored. Some other promising applications include text, where the advent of Large Language Models (LLMs) has facilitated dynamic conversations in various settings like chatbots and healthcare assistants (Gilbert et al., 2023). Studies have indicated that model outputs can be influenced through iterative feedback (Zamfirescu-Pereira et al., 2023; Logan IV et al., 2021), raising fairness concerns. This, coupled with contextual interactions (Zhang et al., 2021), underscores the need to develop evaluation methods capable of assessing long-term fairness to capture these evolving dynamics.
@@ 6.1 Performative Prediction
Performative prediction refers to the dynamic influence of model decisions on future data distribution. At its core is the concept of a distribution map $D:\Theta\rightarrow\mathcal{P}(X\times Y)$, which captures the dependence of data-generating distribution on the model parameters $\theta\in\Theta$, ignoring other reasons for distribution shift (Hardt & Mendler-Dünner, 2023). This abstraction facilitates the modeling of any mechanism for the distribution shift. Miller et al. (2021) propose a two-stage approach where the learner first estimates a model of the distribution map to capture the unknown parameters of the performative effects and then uses the estimated model to construct the proxy of perturbed performative loss. Next, Mendler-Dünner et al. (2020) categorize the stochastic optimization in performative prediction into greedy deploy and lazy deploy. Greedy deployment involves a sequential program that continually updates model parameters based on current data observations and deploys the updated model. In contrast, lazy deployment progressively approximates repeated risk minimization, collecting multiple rounds of observations to update the model between deployments. The survey by Hardt & Mendler-Dünner (2023) offers a comprehensive review of approaches, which can be extended to encompass long-term fairness settings.
@@ 6.2 Delayed Impact
Delayed impact represents how the model decisions under fairness interventions may affect the long-term well-being of the groups it aims to protect for Liu et al. (2018). It can be seen as a case of performative prediction without adopting the formal framework. A couple of works have been brought up to mitigate the delayed impact in the context of multi-armed bandits Tang et al. (2020) or RL Weber et al. (2022). The proposed algorithms include two common steps to mitigate the delayed impact. In the first step, the model learns from the historical data through UCB or optimization as a classification problem. In the second step, the delayed impact will be estimated through importance sampling on the realized future data. The estimated delayed impact will be used as a reward signal that can be further used to select the arms or determine the classifier.
@@ 6.3 Distribution Shift
An alternative research avenue explores dynamic learning environments where data distributions evolve over time. For instance, the profile of loan applicants may shift due to macroeconomic trends or changes in self-selection criteria. This dynamic necessitates continuous fairness maintenance rather than a one-time requirement. A representative work Almuzaini et al. (2022) proposes a dynamic learning method that proactively corrects the classifier, estimating future data properties and retraining accordingly without direct model-environment interaction. Within this framework, two approaches emerge 1) Adaptive, adapting the classifier to a known target environment, and 2) Robust, training classifiers to ensure fairness under perturbations when no target environment information is available (Barrainkua et al., 2023). This problem setting differs from other works due to the absence of direct environmental feedback. For an in-depth review of works in this area, we refer readers to the survey on fairness under domain shift Barrainkua et al. (2023).

### v3

@@ 4.3 Discussion
Compared with the standard static fairness interventions, the summarized mitigation methods mainly contribute to formulating the dynamics of human-model interactions and establishing the corresponding solutions based on the group utility shifts. Long-term fairness under unknown dynamics and/or without leveraging demographic information remains largely unexplored but promising, especially considering privacy laws (Chen et al., 2019a) that limit access to such data. From the perspective of temporal learning, it is positive that long-term fairness is focused on the multi-step sequential learning setting. However, the current research typically only focuses on optimizing the model learner at each time step. Understanding the tensions between ensuring fairness at each time step and cumulatively at the final step needs more investigation. Finally, while RL methods are flexible and learn the dynamics of the environment well, they are data-intensive (Dulac-Arnold et al., 2021; Henderson et al., 2018) and often unsuitable for real-world scenarios. Hence, developing techniques that can efficiently estimate the dynamics is crucial. Notably, the recent advance in performative prediction (Perdomo et al., 2020) provides a new perspective on off-policy learning for modeling the interplay between the model policy and agents. Finally, most current approaches do not account for individuals or groups strategically adapting their behavior in response to deployed algorithms (Estornell et al., 2023). Developing mitigation strategies that are robust to such strategic behavior remains an open challenge and is key to ensuring that fairness interventions remain effective over time.
@@ 5 Evaluation of Long-Term Fairness
Practical applications offer a unique context for long-term fairness considerations, depending on which one would expect context-dependent interpretations of technical findings. Therefore, in this section, we present application-specific fairness inquiries motivated by empirical scenarios.
@@ 5.1 Application-Specific Studies
Liu et al. (2018) focus on lending practices to quantify delayed impacts of decisions (selections) made by institutes (e.g., the bank) on different groups. They represent each group through a score distribution, where higher scores indicate a greater likelihood of positive outcomes. Assuming access to a function that provides the expected score change for individuals, they introduce a one-step feedback model and construct outcome curves, which describe the relation between the expected difference in the mean score after the one-step feedback and the natural parameter regimes (e.g., selection rate of the policy). They found that both Demographic Parity and Equal Opportunity can lead to all possible outcomes, i.e., long-term improvement, stagnation, and decline, in the natural parameter regimes. This study emphasizes the impossibility of foreseeing a fairness criterion’s impact without carefully modeling the delayed outcome.
Labor Market.
@@
The seminal work of Hu & Chen (2018) considers a two-stage dynamic reputational labor market model to capture the long-term and reinforcing nature of disparate outcomes. They propose imposing demographic parity within a temporary labor market to achieve long-term equilibrium in the permanent reputational labor market based on affirmative action. Separately, Celis et al. (2021) investigate the long-term impact of the Rooney Rule (Collins, 2007), a hiring policy aimed at mitigating implicit bias by mandating decision-makers to include at least one candidate from an underrepresented group. They employ a mathematical model of implicit bias, featuring two candidate groups, one of which is underrepresented. Each candidate possesses both a true, latent utility and an observed utility, subject to bias by a multiplicative factor. Candidates are then shortlisted by a selection panel based on observed utility. Their findings demonstrate that the rule progressively enhances minority hiring while maintaining overall utility.
Admission-Hiring Pipeline.
Kannan et al. (2019) study a two-stage model where a student sequentially (1) gets admitted to a college, and then, if admitted, (2) gets hired by an institute. Economic literature has explored combating discrimination through affirmative actions in similar two-stage models, where students invest in human capital (at some cost) during the first stage and enter the labor market in the second stage (Foster & Vohra, 1992; Coate & Loury, 1993). The authors argue that affirmative action interventions have slow effects and advocate studying “downstream effects” for short-term impacts. They consider interventions by colleges in admission rules, grading policies, and information sharing with hiring institutes, and eliminating downstream bias in hiring decisions based on demographic group identity.
Predictive Policing.
@@
Ensign et al. (2018) explore resource allocation in predictive policing (Mohler et al., 2015), which deploys patrol officers based on historical crime data. They identify a feedback phenomenon present in similar algorithms across contexts like recidivism prediction and lending, where prediction outcomes heavily influence algorithm feedback. Using the Pólya urn model, they explain this feedback occurrence and propose black-box solutions, such as reweighing input incidents. Elzayn et al. (2019) go one step further by proposing an algorithm which, under simplification assumptions, converges to an optimally fair resource allocation, even in the absence of known stationary ground-truth crime rates, avoiding the runaway feedback loop effect.
@@ 5.2 Long-Term Fairness in Recommender Systems
Long-term fairness considerations naturally align with Recommender Systems due to evolving user interactions and item popularity, impacting system performance (Liu et al., 2021). Consequently, the RL framework is favored for its ability to model interactive dynamics. The works mostly focus on 1) user-side fairness, addressing equitable treatment for users, and 2) item-side fairness, ensuring unbiased item recommendation stemming from popularity bias due to disproportionate exposure (Chen et al., 2023).
@@
Long-term fairness efforts in recommendation systems primarily target item-side fairness, addressing the challenge by modeling the sequential interactions between consumers and recommender systems using an MDP framework. A few works in this direction adopt an Actor-Critic RL algorithm leveraging the policy-based method (actor) with the value-based approach (critic) to learn the policy and value-function simultaneously (Grondman et al., 2012). In this setup, the actor learns an optimal policy through environment exploration, while the critic evaluates actions to determine long-term rewards (Liu et al., 2021). Utilizing this, Liu et al. (2021) propose FairRec that dynamically maintains accuracy and user fairness in the long term by jointly representing the user preferences (at the item and group level) as the states in an MDP model. They use the Weighted-Proportional Fairness concept, allocating resources (recommendations) proportionally to diverse entities’ demands using a weighting factor. This dual-reward system strategically optimizes both accuracy and fairness in the long term. Separately, Ge et al. (2021) study the long-term fairness of item exposure, focusing on group item labels. They address the challenge of items categorized into groups based on popularity, which dynamically change with user interaction over time, leading to shifting group labels. Using a Demographic Parity-based fairness-constrained policy optimization method, the authors propose an MDP recommendation model based on the actor-critic architecture. However, they introduce an additional critic designed for the demographic parity fairness constraint. On a related note, Morik et al. (2020) explored the rich-get-richer phenomenon in dynamic Learning-to-Rank approaches (Yang & Ai, 2021). Their solution, FairCo, ensures meritocratic fairness by allocating exposure to items proportional to their average relevance. FairCo integrates user feedback and dynamically adjusts its ranking based on the fairness constraint, employing a P-Controller approach for modeling, a popular feedback control mechanism (Bequette, 2003). In another work, Saito et al. (2020) tackles self-selection bias (Chen et al., 2023), where popular items tend to receive more clicks regardless of actual interest. They propose an ideal loss function optimized for relevance rather than the highest click probability. Similarly, Xu et al. (2023a) proposes an online ranking model based on the Max-Min notion of fairness in the batched bandit settings. It balances the fairness-accuracy trade-off by adding a fairness constraint and utilizing the Upper Confidence Bound (UCB) algorithm. In this regard, uniform data has also been proven to be an effective approach (Akpinar et al., 2022). This method involves randomly selecting items and ranking them using a uniform distribution. Chen et al. (2023) (Section 4.8) summarizes the various methods that use this approach. Lastly, Ferraro et al. (2021) investigates gender fairness in collaborative filtering, revealing disparities in exposure between female and male artists. They introduce an iterative re-ranking method that simulates feedback loops and penalizes male artists, resulting in adjusted rankings without significant performance impact.
@@
Relatively fewer works have focused on long-term user-side fairness. Akpinar et al. (2022) employ a simulation framework and Pólya urns to study the long-term impacts of fairness interventions in social network connection recommendations. They argue that when used in dynamic ranking settings, typical statistical notions of fairness fail to maintain equity in network sizes across different protected groups over time. This leads to minority groups having smaller network sizes and lower utility.
@@ 5.3 Discussion
The majority of the long-term fairness evaluations have either been context-specific (e.g., hiring) or in recommendation systems. In recommendation systems, RL methods are favored for modeling dynamics. However, a notable drawback of RL methods is their vulnerability to user tampering (Kasirzadeh & Evans, 2023), where recommender systems manipulate user opinions and preferences to boost long-term user engagement through their recommendations. Furthermore, assessing long-term fairness in a multi-stakeholder (Abdollahpouri & Burke, 2019) setting characterized by competing interests remains largely unexplored. Some other promising applications that have largely been underexplored include text, where the advent of Large Language Models (LLMs) has facilitated dynamic conversations in various settings like chatbots and healthcare assistants (Gilbert et al., 2023). Studies have indicated that model outputs can be influenced through iterative feedback (Zamfirescu-Pereira et al., 2023; Logan IV et al., 2021), raising fairness concerns. This, coupled with contextual interactions (Zhang et al., 2021), underscores the need to develop evaluation methods capable of assessing long-term fairness to capture these evolving dynamics. For instance, incorporating long-term fairness in Reinforcement Learning from Human Feedback (RLHF) (Chaudhari et al., 2024), a popular LLM optimization method, or prompt engineering to debias and capture diverse group preferences is an interesting direction.
@@ 6.1 Performative Prediction
Performative prediction is a recently formalized framework that captures how deploying a predictive model can causally alter the environment or population it operates on, resulting in a model-induced distribution shift (Hardt & Mendler-Dünner, 2023). A key distinction of performative prediction is that it is often formulated as stateless: the environment’s response at each step depends directly on the model’s current action, without explicit dependence on the sequence of past states. This contrasts with dynamic modeling, which explicitly tracks how the state of a system evolves over time based on both past actions and feedback loops. At its core is the concept of a distribution map $D:\Theta\rightarrow\mathcal{P}(X\times Y)$, which captures the dependence of data-generating distribution on the model parameters $\theta\in\Theta$, ignoring other reasons for distribution shift (Mishler & Dalmasso, 2022). This abstraction enables modeling a wide range of mechanisms behind distribution shifts induced by model deployment. Miller et al. (2021) propose a two-stage approach where the learner first estimates a model of the distribution map to capture the unknown parameters of the performative effects and then uses the estimated model to construct the proxy of perturbed performative loss. Next, Mendler-Dünner et al. (2020) categorize the stochastic optimization in performative prediction into greedy deploy and lazy deploy. Greedy deployment involves a sequential program that continually updates model parameters based on current data observations and deploys the updated model. In contrast, lazy deployment progressively approximates repeated risk minimization, collecting multiple rounds of observations to update the model between deployments. The abstract nature of the performative prediction framework also readily lends itself to supporting settings like strategic classification (Perdomo et al., 2020). The survey by Hardt & Mendler-Dünner (2023) offers a review of approaches, which can be extended in future work to encompass long-term fairness settings.
@@ 6.2 Delayed Impact
Delayed impact describes scenarios where the effects of a model or intervention appear over time, not immediately, which can turn short-term gains like bank loan approvals into long-term harms for some groups (Liu et al., 2018). It can be seen as a case of performative prediction without adopting the formal framework (Hardt & Mendler-Dünner, 2023). Importantly, delayed impact differs from performative prediction and dynamic modeling in both scope and modeling assumptions. It typically relies on simpler feedback models like one-step transitions and does not require the ongoing, high-dimensional feedback loops that characterize the other frameworks, which further simplifies both theoretical and empirical analysis. A couple of works have been brought up to mitigate the delayed impact in the context of multi-armed bandits (Tang et al., 2020) and RL (Weber et al., 2022). The proposed algorithms include two common steps to mitigate the delayed impact. In the first step, the model learns from the historical data through UCB or optimization as a classification problem. In the second step, the delayed impact will be estimated through importance sampling on the realized future data. The estimated delayed impact will be used as a reward signal that can be further used to select the arms or determine the classifier. Delayed impact is best suited for evaluating the downstream consequences of a single intervention or a limited set of interventions in relatively stable environments like college admissions and the labor market, as discussed in Section 5.1
@@ 6.3 Distribution Shift
An alternative research avenue explores dynamic learning environments where data distributions evolve over time. For instance, the profile of loan applicants may shift due to macroeconomic trends or changes in self-selection criteria. This dynamic necessitates continuous fairness maintenance rather than a one-time requirement. A representative work proposes a dynamic learning method that proactively corrects the classifier, estimating future data properties and retraining accordingly without direct model-environment interaction (Almuzaini et al., 2022). Within this framework, two approaches emerge 1) Adaptive, adapting the classifier to a known target environment, and 2) Robust, training classifiers to ensure fairness under perturbations when no target environment information is available (Barrainkua et al., 2023). This problem setting differs from other works due to the absence of direct environmental feedback. For an in-depth review of works in this area, we refer readers to the survey on fairness under domain shift (Barrainkua et al., 2023).
7 Critical Analyses and Practical Considerations
In this section, we synthesize insights from existing research to critically assess methodological assumptions and limitations and inform practical considerations.
Questioning Normative Assumptions.
Many fairness approaches implicitly assume that “positive” decisions (e.g., loan approvals, admissions) are inherently beneficial, while “negative” decisions cause harm (Binns, 2018; Chouldechova & Roth, 2018). These underlying normative assumptions, often implicit, shape what outcomes are desirable and how fairness is operationalized in automated decisions (Cooper, 2020). However, this assumption oversimplifies complex realities, e.g., a loan approval does not guarantee a positive outcome if it leads to financial hardship. Existing long-term fairness frameworks, such as effort-based fairness notions (see Section 3.1), do not account for this nuance and tend to conflate a “positive decision” by the algorithm with a positive outcome. Moreover, long-term fairness cannot be meaningfully assessed independently of how predictions are used in policy. The social consequences of algorithmic decisions depend heavily on how those decisions are used in practice, e.g., in credit lending decisions, some policies may aim to extend loans proactively to underserved groups to promote financial inclusion, while others may adopt a conservative approach that restricts credit access to minimize defaults, which might be potentially a better long-term outcome. Zezulka & Genin (2024) demonstrate this dynamic in labor markets, showing that enforcing statistical parity and equal opportunity in risk predictions can undermine efforts to lower overall long-term unemployment and close gender gaps. These findings underscore the need to evaluate fairness through a long-term lens that accounts for downstream consequences, policy interactions, and evolving social dynamics rather than simply extending static fairness metrics or assuming that positive decisions are inherently beneficial. Therefore, discussions of long-term fairness should explicitly consider how positive (and negative) decisions distribute benefits and burdens across groups and individuals over time.
Fragmented Landscape of Long-Term Fairness Notions.
The majority of the static fairness notions, such as statistical parity and equal opportunity, are broadly applicable across diverse domains, resulting in a unified framework for fairness evaluation. This also facilitates the development of standardized benchmarks and comparative studies across models and datasets (Ganesh et al., 2024). In contrast, long-term fairness notions are distinctly fragmented, some are even mutually exclusive, shaped by their close association with specific application domains and the diverse modeling frameworks they rely upon. For example, effort-based fairness, which emphasizes individual progress and cumulative effort, is typically applied outside of the RL framework, making it suitable for domains like education and employment where sequential decision processes or environment feedback are not explicitly modeled. This fragmentation means that these fairness notions and methods are rarely directly comparable or interchangeable, as each emerges from distinct assumptions about temporal dynamics and normative goals inherent to their respective applications and modeling paradigms. This hinders investigation of joint feasibility, or trade-offs between different long-term fairness notions, unlike static fairness literature (Friedler et al., 2019; Bao et al., 2021)
Feasibility of the Causal Modeling of Dynamics.
In terms of principled modeling of underlying dynamics, causal models provide flexible yet precise ways to incorporate understandings of data-generating processes (D’Amour et al., 2020; Creager et al., 2020; Zhang et al., 2020; Zhang & Liu, 2021; Tang et al., 2023a). Existing works tend to (reductively) focus on observed individual-level features when capturing interplays between dynamics, but the complexity of real-world scenarios often calls for a more fine-grained description of data generating processes, potentially among both observable variables (since they are easily measured and readily available) and important latent causal factors (variables that play essential roles in practical scenario but are not particularly easy to measure, e.g., socio-economic status). Tang et al. (2023a) introduce a latent causal factor that carries over current state of affairs to future time steps when defining a long-term fairness goal, and there remains room to explore the role of additional latent factors not yet captured by the current model. Recent developments in causal discovery and causal representation learning have provided theoretical guarantees of the existence and relationship among latent factors, under mild assumptions (Xie et al., 2020; Schölkopf et al., 2021; Huang et al., 2022; Dong et al., 2024; Kong et al., 2024; Zhang et al., 2024; Zheng et al., 2025). The integration of findings from causal representation learning into current modeling approaches remains limited in the current literature on long-term fairness.
The Gap Between Theory and Practice in Long-Term Fairness.
The current body of research on long-term fairness has established rigorous theoretical foundations, particularly through causal inference and reinforcement learning frameworks that model fairness dynamics across temporal and feedback-driven contexts (Creager et al., 2020; Tang et al., 2023a; Yin et al., 2023; Wen et al., 2021). However, practical application remains constrained by limited and simplified empirical problem settings. This is largely due to the scarcity of longitudinal datasets annotated for fairness-relevant outcomes, challenges in capturing real-world environment dynamics (Henderson et al., 2018), and difficulties in managing confounding factors inherent in observational data (Zhang et al., 2020). As a result, empirical experiments often rely on simplistic simulations, synthetic data, and retrospective analyses, which, while useful, do not yet provide a comprehensive assessment or real-world deployment (Le Quy et al., 2022). For example, the foundational benchmark by D’Amour et al. (2020) is currently the only established benchmark to simulate dynamical systems, but it is limited in scope and based on simplified scenarios. Thus, bridging theory and practice requires the development of richer longitudinal datasets and more sophisticated benchmarks capable of modeling complex, real-world dynamics. Recent advances in large reasoning models offer a promising direction toward this goal.